How do I write equation for this one.... The second of two numbers is 4 more than the first. If the second is increased by 1, the result is twice the first. Find the numbers.
Answer by funmath(2873)

Let the first number=x
Then the second number=x+4 (4 more than the 1st)
Increase the second (x+4)by 1 (+1) is (=) twice (2*) the first (x).
(x+4)+1=2x
x+5=2x
-x+x+5=2x-x
5=x
The 1st number:x=5
The 2nd number:x+4=5+4=9
-----------------------------------------------------------------
Check:
9+1=2(5)
10=10
Looks like we're right!
Happy calculating!!!